If n=30, ¯ x x¯ (x-bar)=47, and s=4, find the margin of error at a 99% confidence level
Answer #1

Solution Given that S = n = 30 df - D-I = 30 -1 = 29 A+ 99% confidence lever the tiis x = 1 - 99%. = 1- 0.99 = 0.01 x o.o 0.0